OpenStudy (anonymous):

A cylindrical water trough has a diameter of 6 ft and a height of 2 feet. It is being filled at the rate of 4 ft^3/min, so how fast is the water level rising when the water is 1 foot deep?

OpenStudy (anonymous):

the radius is 6^2 pi feet. it's filling with 4 ft^3/min \[\large \frac{ 4ft^3 /\min}{ 6^2 \pi ft^2}=0.03537 ft/\min\]

OpenStudy (anonymous):

oh wait i messed up the area

OpenStudy (anonymous):

\[\large \frac{ 4ft^3 /\min}{ 3^2 \pi ft^2}=0.1415 ft/\min\]

OpenStudy (anonymous):

i need fraction form :)

OpenStudy (anonymous):

\[\large \frac{ 4 }{ 9 \pi }\]

OpenStudy (anonymous):

thank you so much and i have another question like it ok?